\begin{thebibliography}{18}

\bibitem{tran2009}
M. Tran \textit{et al.}, Phys. Rev. Lett. {\bf 102}, 036601 (2009).

\bibitem{ieeefert2007}
A. Fert \textit{et al.}, IEEE Trans. Electron Devices {\bf 54}, 921 (2007).

\bibitem{valetfert1993}
T. Valet and A. Fert, Phys. Rev. {\bf B48}, 7099 (1993).

\bibitem{fertjaffres2001}
A. Fert and H. Jaffr\`es, Phys. Rev. {\bf B64}, 184420 (2001).

\bibitem{note1}
We remind that $r_1$ writes simply $r_1=\rho~l_{sf}$ in the limit where $l_{sf}\ll W$ and $l_{sf}\ll w$ corresponding to the conventional current-perpendicular-to plane (CPP) geometry (see Ref.~\cite{valetfert1993}).

\bibitem{note2}
For amorphous Co/Al$_2$O$_3$ magnetic tunnel junctions, $\gamma$ is generally taken to be equal to $\gamma=0.4$ like measured by spin-resolved electroluminescence experiments [See L. Lombez \textit{et al}, Appl. Phys. Lett. \textbf{90}, 081111 (2007)].

\bibitem{dzhioev2002}
R.~I. Dzhioev \textit{et al.}, Phys. Rev. {\bf B66}, 245204 (2002).

\bibitem{dash2011}
S.~P. Dash \textit{et al.}, Phys. Rev. {\bf B84}, 054410 (2011).

\end{thebibliography}

\documentclass[prl,showpacs,superscriptaddress,twocolumn]{revtex4-1}
\usepackage{graphicx}		
\usepackage{dcolumn}		
\usepackage{bm}					
\usepackage{textcomp}		

\begin{document}
\title{Erratum : Enhancement of the Spin Accumulation at the Interface Between a Magnetic Tunnel Junction and a Semiconductor.\\ {[}\textbf{Phys. Rev. Lett.~\textbf{102}, 036601 (2009)}]}
\author{M. Tran}
\author{H. Jaffr\`es}
\author{J.-M. George} \email{jean-marie.george@thalesgroup.com}
\author{A. Fert}
\author{A. Miard}
\author{A. Lema\^{\i}tre}

\pacs{72.25.Dc, 75.47.-m, 85.75.-d}
\date{\today}
\maketitle

In the beginning of a previous letter~\cite{tran2009}, we have given a litteral expression for a characteristic threshold resistance $r_1$ of a semiconductor~\cite{ieeefert2007} at an interface with a magnetic tunnel junction as $r_1=(\rho l_{sf})~W/w$ where $\rho$ is the semiconductor resistivity, $l_{sf}$ its spin diffusion length, $W$ the contact width and $w$ the channel thickness. This was derived in the standard theory of spin diffusion/relaxation in the conduction band~\cite{valetfert1993} taking into account a geometric renormalization when the spin transport is made along a lateral channel~\cite{fertjaffres2001}. However, the previous expression is correct only in the limit where $l_{sf}\gg (W,~w)$. In the case of a spin diffusion length well shorter than the contact width $W$ ($l_{sf}\ll W$) but still in the limit $l_{sf}\gg w$ like in the experimental situation, $r_1$ should instead write $r_1=\rho~(l_{sf})^2/w$~\cite{note1}. In this limit, it results that the correct spin resistance area product writes $R_S.A=\gamma^2 r_1=\gamma^2 \rho l_{sf}^2/w\simeq \rho l_{sf}^2/w \simeq 200~\Omega.~\mu m^2$ for a maximal value of $\gamma=1$ ($\gamma$ is the tunnel spin asymmetry coefficient~\cite{note2}) and a maximum value of $l_{sf}=1~\mu$m for the range of doping used~\cite{dzhioev2002}. Since the expected $R_S.A$ product appears even more smaller than the one previously reported (1 $k\Omega.~\mu m^2$), it does not change the core of the present letter emphasizing on the role of the surface states (or localized states) at the direct oxide/semiconductor interfaces giving rise to a strong amplification of the spin signal at the level of 3.6~$M\Omega.~\mu m^2$ or beyond~\cite{dash2011} as experimentally observed.

\end {document}